OpenStudy (anonymous):

Show whether Rolle's theorem can be applied to f(x). If it can, determine any values of c in the interval [0,2pi] for which f'(c)= 0. f(x)= cosx

OpenStudy (anonymous):

so for Rolle's th. to be applicable, f must be continuous and differentiable. I believe this function is. What else needs to be proven?

OpenStudy (anonymous):

I could use some help. @turtlescarf, do you have any ideas?

OpenStudy (anonymous):

c=0 and 2pi

OpenStudy (anonymous):

cosx is differentiable and continuous. f'(c)=sinc. then set sinc=0 and solve for c. you get 0 and 2pi :)

OpenStudy (anonymous):

oh and f(0) must equal f(2pi)

OpenStudy (anonymous):

sorry let me do the problem again, i was using medium value theorem

OpenStudy (anonymous):

okay, cos0 and cos2pi both equal 1. therefore, f'(c)=0

OpenStudy (anonymous):

oh thank you for the responses! I haven't read them yet so give me a second.

OpenStudy (anonymous):

so I just need to prove that cos(0)= cos(2pi)? along with the fact that f is continuous and differentiable. And that would prove Rolle's Th. Applicable?

OpenStudy (anonymous):

oh but then what would c be? how do I solve for that?

OpenStudy (anonymous):

find the derivative and set = to 0 and solve for x?

OpenStudy (anonymous):

@turtlescarf ?

OpenStudy (anonymous):

set f'(c)=0 so you can solve for c

OpenStudy (anonymous):

you're right, but you solve for c instead of x

OpenStudy (anonymous):

so if f'(x)= -sinx and I set equal to 0, how do I solve with the 'sin' wait how do I solve for the c?

OpenStudy (anonymous):

f'(c)=-sinc

OpenStudy (anonymous):

you replace the x with c

OpenStudy (anonymous):

then, using the theorem, you set -sinc=0

OpenStudy (anonymous):

and then set to 0?

OpenStudy (anonymous):

divide both sides by -1 and you get sinc=0

OpenStudy (anonymous):

so c is 0 and 2pi

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

oh ok so c could be either because they are both in the interval?

OpenStudy (anonymous):

0 and 2pi are technically the same thing on the circle*

OpenStudy (anonymous):

hm i'd say c is both, even though both values are the same, just in case your teacher/whoever prefers a certain answer

OpenStudy (anonymous):

basically, yes.

OpenStudy (anonymous):

oh ok. Thank you so much for your help! I'll give you a medal and then close this question because I have one more problem to get to tonight!

OpenStudy (anonymous):

in response to "so I just need to prove that cos(0)= cos(2pi)? along with the fact that f is continuous and differentiable. And that would prove Rolle's Th. Applicable?" Yes. and also show that f(a)=f(b), which in this case is f(0)-f(2pi).

OpenStudy (anonymous):

you're welcome